# Detecting Change Points of Covariance Matrices in High Dimensions

### v1

@@ Abstract
Testing for change points in sequences of high-dimensional covariance matrices is an important and equally challenging problem in statistical methodology with applications in various fields. Motivated by the observation that even in cases where the ratio between dimension and sample size is as small as $0.05$, tests based on a fixed-dimension asymptotics do not keep their preassigned level, we propose to derive critical values of test statistics using an asymptotic regime where the dimension diverges at the same rate as the sample size.
This paper introduces a novel and well-founded statistical methodology for detecting change points in a sequence of high-dimensional covariance matrices. Our approach utilizes a min-type statistic based on a sequential process of likelihood ratio statistics. This is used to construct a test for the hypothesis of the existence of a change point with a corresponding estimator for its location. We provide theoretical guarantees for these inference tools by thoroughly analyzing the asymptotic properties of the sequential process of likelihood ratio statistics in the case where the dimension and sample size converge with the same rate to infinity. In particular, we prove weak convergence towards a Gaussian process under the null hypothesis of no change. To identify the challenging dependency structure between consecutive test statistics, we employ tools from random matrix theory and stochastic processes. Moreover, we show that the new test attains power under a class of alternatives reflecting changes in the bulk of the spectrum, and we prove consistency of the estimator for the change-point location.
15A18,
60F17,
62H15,
high-dimensional change point analysis,
likelihood ratio test,
covariance matrices,
substitution principle,
@@ 1 Introduction
Having its origins in quality control (see Wald, 1945; Page, 1954, for two early references), change point detection has been an extremely active field of research until today with numerous applications in finance, genetics, seismology or sports to name just a few. In the last decade, a large part of the literature on change point detection considers the problem of detecting a change point in a high-dimensional sequence of means (see Jirak, 2015; Cho and Fryzlewicz, 2015; Dette and Gösmann, 2020; Enikeeva and Harchaoui, 2019; Liu et al., 2020; Liu, Gao and Samworth, 2021; Chen, Wang and Wu, 2022; Wang et al., 2022; Zhang, Wang and Shao, 2022, among many others).
Compared to the huge amount of work on the change-point problem for a sequence of high-dimensional means, the literature on the problem of detecting structural breaks in the corresponding covariance matrices is more scarce. For the low dimensional setting we refer to Chen and Gupta (2004), Lavielle and Teyssiere (2006), Galeano and Peña (2007), Aue et al. (2009) and Dette and Wied (2016), among others, who study different methods and aspects of the change point problem under the assumption that the sample size converges to infinity while the dimension is fixed. We also refer to Theorem 1.1.2 in Csörgő and Horváth (1997) who provide a test statistic and its asymptotic distribution under the null hypothesis for normal distributed data. However, even in cases where the ratio between dimension and sample size is rather small, it can be observed that statistical guarantees derived from fixed-dimension asymptotics can me be misleading. Exemplary, we display in Table 1 the simulated type I error of two commonly used tests for a change point in the in a sequence of covariance matrices. The first method (CH) is based on sequential likelihood ratio statistics, where the critical values have been determined by classical asymptotic arguments assuming that the dimension is fixed (see Theorem 1.1.2 in Csörgő and Horváth, 1997). The second approach (AHHR) is a test proposed by Aue et al. (2009), which is based on a quadratic form of the vectorized CUSUM statistic of the empirical covariance matrix. Again, the determination of critical values relies on fixed-dimensional asymptotics. We observe that even in the case where the ratio between the dimension and sample size is as small as $0.05$, the nominal level $\alpha=0.05$ of the CH test is exceeded by more than a factor of three. On the other hand, the AHHR test provides only a reasonable approximation of the nominal level, if the ratio between dimension and sample size is $0.025$. Note that this test requires the inversion of an estimate of high-dimensional covariance matrix and is only applicable if the sample size is larger than the squared dimension.
@@
Table 1: Simulated type I errors of the sequential likelihood ratio test (Theorem 1.1.2 in Csörgő and Horváth, 1997) and the test of Aue et al. (2009) ($500$ simulation runs, nominal level $\alpha=0.05$, standard normally distributed data). Critical values are determined by fixed dimension asymptotics for a sample size of $n=200$. If ”-” is reported, the corresponding test is not applicable.
@@
Meanwhile, several authors have also discussed the problem of estimating a change point in a sequence of covariance matrices in the high-dimensional regime. For example, Avanesov and Buzun (2018) propose a multiscale approach to estimate multiple change points, while Wang, Yu and Rinaldo (2021) investigate the optimality of binary and wild binary segmentation for multiple change point detection. We also mention the work of Dette, Pan and Yang (2022), who propose a two-stage approach to detect the location of a change point in a sequence of very high-dimensional covariance matrices. Li and Gao (2024) pursue a similar approach to develop a change-point test for high-dimensional correlation matrices.
@@
With respect to testing, the literature on change point detection is rather scarce. In principle, one can develop change point analysis based on a vectorization of the covariance matrices using inference tools for a sequence of means. This approach essentially boils down to comparing the matrices before and after the change point with respect to a vector norm. However, in general, this approach does not yield an asymptotically distribution free test statistic. Moreover, as pointed out by Ryan and Killick (2023), such distances do not reflect the geometry induced on the space of positive definite matrices. These authors propose a change-point test based on an alternative distance defined on the space of positive definite matrices, which compares sequentially the multivariate ratio $\mathbf{\Sigma}^{-1}_{1}\mathbf{\Sigma}_{2}$ of the two covariance matrices $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{2}$ before and after a potential change point with the identity matrix. As a consequence, under the null hypothesis of no change point, their test statistic is independent of the underlying covariance structure, which makes it possible to derive quantiles for statistical testing in the regime where the dimension diverges at the same rate as the sample size. However, the approach of these authors is based on a combination of a point-wise limit theorem from random matrix theory with a Bonferroni correction. Therefore, as pointed out in Section 4 of Ryan and Killick (2023), the resulting test may be conservative in applications. Moreover, this methodology is tailored to centered data, and it is demonstrated in Zheng, Bai and Yao (2015), that an empirical centering introduces a non-negligible bias in the central limit theorem for the corresponding linear spectral statistic.
@@
In this paper, we propose an alternative test for detecting a change point in a sequence of high-dimension covariance matrices, which takes the strong dependence between consecutive test statistics into account to avoid the drawbacks of previous works. Our approach is based on a sequential process of likelihood ratio test (LRT) statistics, where the dimension of the data is growing with the dimension at the same rate. We combine tools from random matrix theory and stochastic processes to develop and analyze statistical methodology for change point analysis in the covariance structure of high-dimensional data. Random matrix theory is a common tool to investigate asymptotic properties of LRT in high-dimensional scenarios for classical testing problems. An early reference in this direction is Jiang and Yang (2013), who establish central limit theorems for several classical LRT statistics under the null hypotheses. Since this seminal work, numerous researchers have investigated related problems (see Jiang and Qi, 2015; Dette and Dörnemann, 2020; Bao et al., 2022; Dörnemann, 2023; Heiny and Parolya, 2023; Parolya, Heiny and Kurowicka, 2024, among others). None of these papers considers sequential LRT statistics to develop change point analysis. Moreover, these references do not contain any consistency results. In general, the body of literature concerning LRTs under alternative hypotheses in the high-dimensional regime remains sparse; for some exceptions, see Chen and Jiang (2018), Bodnar, Dette and Parolya (2019) and Bai and Zhang (2023). Having this line of literature in mind, we can summarize the main contributions of this paper.
  1. 1.
@@
    Enhancing the power of the min-type statistic. We propose a novel methodology to test for a change point in a sequence of high-dimensional covariance matrices based on a minimum of sequential LRT statistics. Under the null hypothesis, this statistic admits a simple limiting distribution in the regime where the dimension diverges proportionally to the sample size. Unlike most other approaches, the new test does not require the estimation of the population covariance matrix. This result facilities the introduction of a simple asymptotic testing procedure with favorable finite-sample properties. Most notably, our approach takes the strong dependency structure between consecutive test statistics into account, whose analysis has been recognized as a challenging problem in the literature (see Ryan and Killick, 2023), and which has not been addressed in previous works.
  2. 2.
    Power analysis and change-point estimation. We also study asymptotic properties of the sequential process under alternative hypotheses to identify conditions which guarantee the consistency of the test. The techniques developed in this paper are believed to be a valuable contribution to study consistency for other problems as well.
    Besides developing a testing methodology with desirable properties, we also introduce a new estimator of the change-point location based on the LRT principle. Our comprehensive analysis of the process of LRT statistics enables us to provide theoretical guarantees under which the estimator admits consistency.
  3. 3.
@@
    Mathematical challenges. Investigating sequential statistics introduces new challenges in random matrix theory compared to consideration of the common (non-sequential) LRT, namely (i) the convergence of the finite-dimensional distributions and (ii) the asymptotic tightness of the sequential log-LRT statistics. Indeed, the weak convergence result implied by (i) and (ii) is a novel, technically challenging contribution, given that sequential LRT statistics have not been studied in such a functional and high-dimensional framework before.
@@
    To establish (i), we derive an asymptotic representation of the test statistics and apply a martingale CLT to the dominating term in this decomposition. Note that for given time points $t_{1},t_{2}\in[0,1]$, the corresponding LRT statistics are highly correlated, and a nuanced analysis is required in order to determine their covariance. Regarding (ii), we show asymptotic equicontinuity of the sequential log-LRT statistics by deriving uniform inequalities for the moments of the corresponding increments.
@@
The remaining part of this work is structured as follows. In Section 2, we present the new method to detect and estimate a change-point in a high-dimensional covariance structure, and provide the main theoretical guarantees. In numerical experiments given in Section 3, we compare the finite-sample size properties of our test as well as the change-point estimator to other approaches. The proofs of our theoretical results are deferred to Section 4 and the supplementary material.
@@ 2 Change point analysis by a sequential LRT process
Let $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$ be a sample of independent random vectors such that $\mathbf{y}_{i}=(y_{1i},\ldots,y_{pi})^{\top}=\mathbf{\Sigma}_{i}^{1/2}\mathbf{x}_{i}$ for i.i.d. $p$-dimensional random vectors $\mathbf{x}_{i}$ and covariance matrices $\mathbf{\Sigma}_{i}=\mathbf{\Sigma}_{i,n}$, $1\leq i\leq n.$ We are interested in testing for a change in the covariance structure of $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$, and consider the hypotheses
$H_{0}:$ $\mathbf{\Sigma}_{1}=\ldots=\mathbf{\Sigma}_{n}$ (2.1)
@@
$H_{1}:$ $\mathbf{\Sigma}_{1}=\ldots=\mathbf{\Sigma}_{\lfloor nt^{\star}\rfloor}\neq\mathbf{\Sigma}_{\lfloor nt^{\star}\rfloor+1}=\ldots=\mathbf{\Sigma}_{n},$ (2.2)
@@
$\hat{\mathbf{\Sigma}}_{i:j}^{\textnormal{cen}}$ $=\frac{1}{j-i}\sum\limits_{k=i}^{j}\left(\mathbf{y}_{j}-\overline{\mathbf{y}}_{i:j}\right)\left(\mathbf{y}_{j}-\overline{\mathbf{y}}_{i:j}\right)^{\top},\quad 1\leq i\leq j\leq n,$ (2.3)
@@
$\overline{\mathbf{y}}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{y}_{i}$
@@
$\hat{\mathbf{\Sigma}}^{\textnormal{cen}}$ $=\hat{\mathbf{\Sigma}}_{1:n}^{\textnormal{cen}},$ (2.4)
@@
$\Lambda_{n,t}^{\textnormal{cen}}=\frac{\big{|}\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\big{|}^{\frac{1}{2}\lfloor nt\rfloor}\big{|}\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\big{|}^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big{|}\hat{\mathbf{\Sigma}}^{\textnormal{cen}}\big{|}^{\frac{1}{2}n}},\quad t\in(0,1).$ (2.5)
If, for fixed $t$, $\mathbf{y}_{1},\ldots,\mathbf{y}_{\lfloor nt\rfloor}$ and $\mathbf{y}_{\lfloor nt\rfloor+1},\ldots,\mathbf{y}_{n}$ are two independent samples of i.i.d. random variables with $\mathbb{E}[\mathbf{y}_{1}]=\mu_{1}$, Var$(\mathbf{y}_{1})=\mathbf{\Sigma}_{1}$ and $\mathbb{E}[\mathbf{y}_{n}]=\mu_{n}$, Var$(\mathbf{y}_{n})=\mathbf{\Sigma}_{n}$, then $\Lambda_{n,t}^{\textnormal{cen}}$ is the likelihood ratio test statistic (LRT) for the hypotheses $\tilde{H}_{0}:~{}\mathbf{\Sigma}_{1}=\mathbf{\Sigma}_{n},~{}\mu_{1}=\mu_{n}$ versus $\tilde{H}_{1}:~{}\mathbf{\Sigma}_{1}\neq\mathbf{\Sigma}_{n}$. This problem has been investigated by several authors in the high-dimensional regime (see, for example, Li and Chen, 2012; Jiang and Yang, 2013; Dörnemann, 2023; Dette and Dörnemann, 2020; Jiang and Qi, 2015; Guo and Qi, 2024). In contrast to these works, consistent change point inference on the basis of likelihood ratio tests requires the analysis of the full process $(\Lambda_{n,t}^{\textnormal{cen}})_{t\in[t_{0},1-t_{0}]}$. In particular, we need an appropriate centering and standardization of $\log\Lambda_{n,t}^{\textnormal{cen}},$ which is discussed next.
An important ingredient for the centering the statistic $\Lambda_{n,t}^{\textnormal{cen}}$ an estimator of the kurtosis
$\kappa=\mathbb{E}[x_{11}^{4}]$
@@
$\kappa=3+\frac{\operatorname{Var}\left(\|\mathbf{y}_{1}-\mathbb{E}[\mathbf{y}_{1}]\|_{2}^{2}\right)-2\|\mathbf{\Sigma}\|_{F}^{2}}{\sum_{j=1}^{p}\Sigma_{jj}^{4}}.$ (2.6)
@@
$\hat{\tau}_{n}$ $=\operatorname{tr}\big{(}(\hat{\mathbf{\Sigma}}^{\textnormal{cen}})^{2}\big{)}-\frac{1}{n}\big{(}\operatorname{tr}\hat{\mathbf{\Sigma}}^{\textnormal{cen}}\big{)}^{2},$
$\hat{\nu}_{n}$ $=\frac{1}{n-1}\sum_{i=1}^{n}\left(\|\mathbf{y}_{i}-\overline{\mathbf{y}}_{i}\|_{2}^{2}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}\|\mathbf{y}_{i^{\prime}}-\overline{\mathbf{y}}_{i^{\prime}}\|_{2}^{2}\right)^{2},$
@@
$\hat{\kappa}_{n}$ $=\max\Big{\{}3+\frac{\hat{\nu}_{n}-2\hat{\tau}_{n}}{\hat{\omega}_{n}},1\Big{\}}$
(note that a similar estimator was proposed by Lopes, Blandino and Aue (2019) for the case $\mathbb{E}[y_{11}]=0$). A consistency result for $\hat{\kappa}_{n}$ under the null hypothesis will be provided in Lemma 9. We will show in Theorem 1 below that under the null hypothesis we can approximate the expected value and the variance of $2\log\Lambda_{n,t}^{\textnormal{cen}}$ by
@@
Under the alternative $H_{1}$, we will show that (under appropriate assumptions) that
$\frac{2\log\Lambda^{(\textnormal{cen})}_{n,t^{\star}}-\tilde{\mu}_{n,t^{\star}}}{n}=E_{n}(t^{*})+\mathcal{O}_{\mathbb{P}}(1),$ (2.9)
where
$E_{n}(t^{*})=\frac{\lfloor nt^{\star}\rfloor}{n}\log\left|\mathbf{\Sigma}_{1}\right|+\frac{n-\lfloor nt^{\star}\rfloor}{n}\log\left|\mathbf{\Sigma}_{n}\right|-\log\left|\frac{\lfloor nt^{\star}\rfloor}{n}\mathbf{\Sigma}_{1}+\frac{n-\lfloor nt^{\star}\rfloor}{n}\mathbf{\Sigma}_{n}\right|.$ (2.10)
By the log-concavity of the mapping $\mathbf{A}\to\log|\mathbf{A}|$ defined on the set of positive definite matrices we have $E_{n}\leq 0$ with equality if and only if $\mathbf{\Sigma}_{1}=\mathbf{\Sigma}_{n}$ (which means that there is no change point). Therefore, we propose to reject the null hypothesis (2.1) in favor of (2.2) for small values of $M_{n}^{\textnormal{cen}}$. To find corresponding quantiles, we first determine the asymptotic distribution of the test statistic $M_{n}^{\textnormal{cen}}$ under the null hypothesis in the high-dimensional regime, where the dimension diverges at the same rate as the sample size. For this purpose, we make the following assumptions.
  1. (A1)
  2. (A2)
  3. (A3)
    We have uniformly with respect to $n\in\mathbb{N}$
    $0<\lambda_{\min}(\mathbf{\Sigma}_{1})\leq\lambda_{\max}(\mathbf{\Sigma}_{1})<\infty.$
@@
If Assumption (A1), (A2) with $\delta>4$ and Assumption (A3) are satisfied, then we have under $H_{0}$

### v2

# Sequential Eigenvalue Statistics for Change-Point Detection in Covariance Matrices

@@ Abstract
Testing for change points in sequences of covariance matrices is an important and equally challenging problem in statistical methodology with applications in various fields. Motivated by the observation that even in cases where the ratio between dimension and sample size is as small as $0.05$, tests based on a fixed-dimension asymptotics do not keep their preassigned level, we propose to derive critical values of test statistics using an asymptotic regime where the dimension diverges at the same rate as the sample size. This paper introduces a novel and well-founded statistical methodology for detecting change points in a sequence of moderately dimensional covariance matrices. Our approach utilizes a min-type statistic based on a sequential process of likelihood ratio statistics. This is used to construct a test for the hypothesis of the existence of a change point with a corresponding estimator for its location. We provide theoretical guarantees by thoroughly analyzing the asymptotic properties of the sequential process of likelihood ratio statistics. In particular, we prove weak convergence towards a Gaussian process under the null hypothesis of no change. To identify the challenging dependency structure between consecutive test statistics, we employ tools from random matrix theory and stochastic processes.
@@ 1 Introduction
Having its origins in quality control (see Wald, 1945; Page, 1954, for two early references), change point detection has been an extremely active field of research until today with numerous applications in finance, genetics, seismology or sports to name just a few. In the last decade, a large part of the literature on change point detection considers the problem of detecting a change point in a high-dimensional sequence of means (see Jirak, 2015; Cho and Fryzlewicz, 2015; Dette and Gösmann, 2020; Enikeeva and Harchaoui, 2019; Liu et al., 2020, 2021; Chen et al., 2022; Wang et al., 2022; Zhang et al., 2022a, among many others).
Compared to the vast body of work on the change-point problem for a sequence of high-dimensional means, the literature on the problem of detecting structural breaks in the corresponding covariance matrices is relatively scarce. For the low dimensional setting we refer to Chen and Gupta (2004), Lavielle and Teyssiere (2006), Galeano and Peña (2007), Aue et al. (2009) and Dette and Wied (2016), among others, who study different methods and aspects of the change point problem under the assumption that the sample size converges to infinity while the dimension is fixed. We also refer to Theorem 1.1.2 in Csörgő and Horváth (1997) who provide a test statistic and its asymptotic distribution under the null hypothesis for normally distributed data. However, even in cases where the ratio between dimension and sample size is rather small, it can be observed that statistical guarantees derived from fixed-dimension asymptotics can be misleading. For instance, we display in Table 1 the simulated type I error of two commonly used tests for a change point in a sequence of covariance matrices. The first method (CH) is based on sequential likelihood ratio statistics, where the critical values have been determined by classical asymptotic arguments assuming that the dimension is fixed (see Theorem 1.1.2 in Csörgő and Horváth, 1997). The second approach (AHHR) is a test proposed by Aue et al. (2009), which is based on a quadratic form of the vectorized CUSUM statistic of the empirical covariance matrix. Again, the determination of critical values relies on fixed-dimensional asymptotics. We observe that even in the case where the ratio between the dimension and sample size is as small as $0.05$, the nominal level $\alpha=0.05$ of the CH test is exceeded by more than a factor of three. On the other hand, the AHHR test provides only a reasonable approximation of the nominal level if the ratio between dimension and sample size is $0.025$. Note that this test requires the inversion of an estimate of a large dimensional covariance matrix and is only applicable if the sample size is larger than the squared dimension.
@@
Table 1: Simulated type I errors of the sequential likelihood ratio test (Theorem 1.1.2 in Csörgő and Horváth, 1997) and the test of Aue et al. (2009) for a sample size of $n=200$ (500500 simulation runs, nominal level $\alpha=0.05$, standard normally distributed data). Critical values are determined by fixed dimension asymptotics. If ”-” is reported, the corresponding test is not applicable.
@@
Meanwhile, several authors have also discussed the problem of estimating a change point in a sequence of covariance matrices in the high-dimensional regime. For example, Avanesov and Buzun (2018) propose a multiscale approach to estimate multiple change points, while Wang et al. (2021) investigate the optimality of binary and wild binary segmentation for multiple change point detection. We further mention the work of Dette et al. (2022), who propose a two-stage approach to detect the location of a change point in a sequence of very high-dimensional covariance matrices. Li and Gao (2024) pursue a similar approach to develop a change-point test for high-dimensional correlation matrices.
@@
The literature on testing for change points is relatively scarce. In principle, one can develop change point analysis based on a vectorization of the covariance matrices using inference tools for a sequence of means. This approach essentially boils down to comparing the matrices before and after the change point with respect to a vector norm. However, in general, this approach does not yield an asymptotically distribution free test statistic. Moreover, as pointed out by Ryan and Killick (2023), such distances do not reflect the geometry induced on the space of positive definite matrices. Their work introduces a change-point test based on an alternative distance defined on the space of positive definite matrices, which compares sequentially the multivariate ratio $\boldsymbol{\Sigma}^{-1}_{1}\boldsymbol{\Sigma}_{2}$ of the two covariance matrices $\boldsymbol{\Sigma}_{1}$ and $\boldsymbol{\Sigma}_{2}$ before and after a potential change point with the identity matrix. As a consequence, under the null hypothesis of no change point, their test statistic is independent of the underlying covariance structure, which makes it possible to derive quantiles for statistical testing in the regime where the dimension diverges at the same rate as the sample size. However, the approach of these authors is based on a combination of a point-wise limit theorem from random matrix theory with a Bonferroni correction. Therefore, as pointed out in Section 4 of Ryan and Killick (2023), the resulting test may be conservative in applications. Moreover, this methodology is tailored to centered data, and it is demonstrated in Zheng et al. (2015), that an empirical centering introduces a non-negligible bias in the central limit theorem for the corresponding linear spectral statistic.
@@
In this paper, we propose an alternative test for detecting a change point in a sequence of covariance matrices, which takes the strong dependence between consecutive test statistics into account to avoid the drawbacks of previous works. Our approach is based on a sequential process of likelihood ratio test (LRT) statistics, where the dimension of the data grows at the same rate as the sample size. We combine tools from random matrix theory and stochastic processes to develop and analyze statistical methodology for change point analysis in the covariance structure. Random matrix theory is a common tool to investigate asymptotic properties of LRT in moderately high-dimensional scenarios for classical testing problems. An early reference in this direction is Bai et al. (2009), who study one- and two-sample problems for covariance matrices and provide Gaussian approximations for LRTs in high dimensions. Moreover, Jiang and Yang (2013) establish central limit theorems for several classical LRT statistics under the null hypotheses. Both works rely on the normal assumption.
Since these seminal works, numerous researchers have investigated related problems (see Jiang and Qi, 2015; Dette and Dörnemann, 2020; Bao et al., 2022; Dörnemann, 2023; Heiny and Parolya, 2024, among others). None of these papers considers sequential LRT statistics to develop change point analysis. Moreover, our approach is conceptually different from most existing work on testing for change points in high-dimensional data (see, for example, Liu et al. (2021) for changes in a mean vector and Wang and Yao (2021) for changes in a covariance matrix) and does neither require a sparsity nor a sub-Gaussian assumption. Having this line of literature in mind, we can summarize the main contributions of this paper.
@@
  • We propose a novel methodology to test for a change point in a sequence of moderately high-dimensional covariance matrices based on a minimum of sequential LRT statistics. Under the null hypothesis, this statistic admits a simple limiting distribution in the regime where the dimension diverges proportionally to the sample size. Unlike most other approaches, the distribution of the test statistic under the null hypothesis is invariant to the population covariance matrix. This result facilitates the introduction of a simple asymptotic testing procedure with favorable finite-sample properties. Most notably, our approach takes the strong dependence structure between consecutive test statistics into account, whose analysis has been recognized as a challenging problem in the literature (see Ryan and Killick, 2023), and which has not been addressed in previous works.
@@
  • Investigating sequential statistics introduces new mathematical challenges compared to the analysis of the standard (non-sequential) LRT, namely (i) the convergence of the finite-dimensional distributions and (ii) the asymptotic tightness of the sequential log-LRT statistics. Indeed, the weak convergence result implied by (i) and (ii) is a novel, technically challenging contribution, given that sequential LRT statistics have not been studied in such a framework before.
@@
    To establish (i), we derive an asymptotic representation of the test statistics and apply a martingale CLT to the dominating term in this decomposition. Note that for given time points $t_{1},t_{2}\in[0,1]$, the corresponding LRT statistics are highly correlated, and a nuanced analysis is required to determine their covariance. Regarding (ii), we show asymptotic equicontinuity of the sequential log-LRT statistics by deriving uniform inequalities for the moments of the increments of the process.
  • Along the way, we develop a consistent estimator of the kurtosis. As numerous results in random matrix theory and high-dimensional statistics demonstrate that spectral statistics depend critically on whether the kurtosis equals three (Bai and Silverstein, 2010; Zhang et al., 2022b; Pan and Zhou, 2008; Zheng, 2012; Yin et al., 2023), this estimator is believed to be of independent methodological interest.
@@
The remaining part of this work is structured as follows. In Section 2, we present the new method to detect a change-point in a covariance structure of moderate dimension, and provide the main theoretical guarantees. In numerical experiments given in Section 3, we compare the finite-sample size properties of our test as well as the change-point estimator to other approaches. The proofs of our theoretical results are deferred to Section 4 and the supplementary material.
@@ 2 Change point analysis by a sequential LRT process
Let $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$ be a sample of independent random vectors such that $\mathbf{y}_{i}=(y_{1i},\ldots,y_{pi})^{\top}=\boldsymbol{\Sigma}_{i}^{1/2}\mathbf{x}_{i}$ for i.i.d. $p$-dimensional random vectors $\mathbf{x}_{i}$ and covariance matrices $\boldsymbol{\Sigma}_{i}=\boldsymbol{\Sigma}_{i,n}$, $1\leq i\leq n.$ We are interested in testing for a change in the covariance structure of $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$, and consider the hypotheses
$H_{0}:$ $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{n}$ (2.1)
@@
$H_{1}:$ $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{\lfloor nt^{\star}\rfloor}\neq\boldsymbol{\Sigma}_{\lfloor nt^{\star}\rfloor+1}=\ldots=\boldsymbol{\Sigma}_{n},$ (2.2)
@@
$\hat{\boldsymbol{\Sigma}}_{i:j}^{\textnormal{cen}}$ $=\frac{1}{j-i}\sum\limits_{k=i}^{j}\left(\mathbf{y}_{k}-\overline{\mathbf{y}}_{i:j}\right)\left(\mathbf{y}_{k}-\overline{\mathbf{y}}_{i:j}\right)^{\top},\quad 1\leq i\leq j\leq n,$ (2.3)
@@
$\overline{\mathbf{y}}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{y}_{k}$
@@
$\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}$ $=\hat{\boldsymbol{\Sigma}}_{1:n}^{\textnormal{cen}},$ (2.4)
@@
$\Lambda_{n,t}^{\textnormal{cen}}=\frac{\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\big|^{\frac{1}{2}\lfloor nt\rfloor}\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\big|^{\frac{1}{2}(n-\lfloor nt\rfloor)}}{\big|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\big|^{\frac{1}{2}n}},\quad t\in(0,1).$ (2.5)
[…]
To formulate the statistical properties of this process, we make the following assumptions.
  1. (A-1)
  2. (A-2)
  3. (A-3)
    We have uniformly with respect to $n\in\mathbb{N}$
    $0<\lambda_{\min}(\boldsymbol{\Sigma}_{1})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{1})<\infty.$
An important ingredient for an appropriate centering of $\log\Lambda_{n,t}^{\textnormal{cen}},$ is an estimator of the kurtosis
$\kappa_{n}=\mathbb{E}[x_{11}^{4}]$
@@
$\kappa_{n}=3+\frac{\operatorname{Var}\left(\|\mathbf{y}_{1}-\mathbb{E}[\mathbf{y}_{1}]\|_{2}^{2}\right)-2\|\boldsymbol{\Sigma}\|_{F}^{2}}{\sum_{j=1}^{p}\Sigma_{jj}^{4}}.$ (2.6)
@@
$\hat{\tau}_{n}$ $=\operatorname{tr}\big((\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}})^{2}\big)-\frac{1}{n}\big(\operatorname{tr}\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\big)^{2},$
$\hat{\nu}_{n}$ $=\frac{1}{n-1}\sum_{i=1}^{n}\left(\|\mathbf{y}_{i}-\overline{\mathbf{y}}\|_{2}^{2}-\frac{1}{n}\sum_{i^{\prime}=1}^{n}\|\mathbf{y}_{i^{\prime}}-\overline{\mathbf{y}}\|_{2}^{2}\right)^{2},$
@@
$\hat{\kappa}_{n}$ $=\max\Big\{3+\frac{\hat{\nu}_{n}-2\hat{\tau}_{n}}{\hat{\omega}_{n}},1\Big\}.$
Our first result provides the consistency of $\hat{\kappa}_{n}$ for $\kappa_{n}$ under the null hypothesis. Its proof is postponed to Section A.4.
Proposition 1.
Suppose that assumptions (A-2)-(A-3) are satisfied, and $p/n\to y\in(0,\infty)$ as $n\to\infty$. Then, under $H_{0}$, we have
$\frac{\hat{\kappa}_{n}}{\kappa_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1.$
Remark 1.
In the case $\mathbb{E}[y_{11}]=0$, a related estimator for $\kappa_{n}$ was proposed by Lopes et al. (2019). To the best of our knowledge, $\hat{\kappa}_{n}$ is the first estimator to be equipped with theoretical guarantees under general (possibly nonzero) means in the regime where the dimension is asymptotically proportionally to the sample size. This estimator is believed to be of independent methodological interest. Indeed, the excess kurtosis $\kappa_{n}-3$ is a key quantity in extending asymptotic results for spectral statistics from the Gaussian case to non-Gaussian settings, see, e.g., Bai and Silverstein (2010); Pan and Zhou (2008); Zheng (2012); Yin et al. (2023); Najim and Yao (2016); Zhang et al. (2022b). Since this parameter directly affects the limiting behavior of eigenvalue-based statistics, we expect that $\hat{\kappa}_{n}$ will find applications in testing problems for moderately high-dimensional data.
We will show that under the null hypothesis we can approximate the expected value and the variance of $2\log\Lambda_{n,t}^{\textnormal{cen}}$ by
@@
In the next theorem, we provide the limiting distribution of $M_{n}^{\textnormal{cen}}$ under the null hypothesis of no change.
@@
If Assumption (A-1), (A-2) with $\delta>4$ and Assumption (A-3) are satisfied, then we have under $H_{0}$